OpenStudy (anonymous):

let f(x)=4+5x and g(x)=2x-1Find f(g(x)) and g(f(x))

OpenStudy (nenadmatematika):

f(g(x))=4+5(2x-1)=10x-1 and g(f(x))=2(4+5x)-1=10x+7